Chamberlain Enterprises Inc. reported the following receivables in its December 31, 2018, year-end balance sheet:

Current assets:
Accounts receivable, net of $24,000 in allowance for
uncollectible accounts
$ 218,000
Interest receivable 6,800
Notes receivable 260,000


Additional Information:

  1. The notes receivable account consists of two notes, a $60,000 note and a $200,000 note. The $60,000 note is dated October 31, 2018, with principal and interest payable on October 31, 2019. The $200,000 note is dated June 30, 2018, with principal and 6% interest payable on June 30, 2019.
  2. During 2019, sales revenue totaled $1,340,000, $1,280,000 cash was collected from customers, and $22,000 in accounts receivable were written off. All sales are made on a credit basis. Bad debt expense is recorded at year-end by adjusting the allowance account to an amount equal to 10% of year-end accounts receivable.
  3. On March 31, 2019, the $200,000 note receivable was discounted at the Bank of Commerce. The bank's discount rate is 8%. Chamberlain accounts for the discounting as a sale.

1. Not including sales revenue, what revenue and expense amounts related to receivables will appear in Chamberlain’s 2019 income statement?
Answer #1
Chamberlain’s income statement will include sales revenue of $1,340,000 plus, interest revenue of $7000, bad debt expense of $26000 and loss on sale of note receivable $1240. The calculations are as follow:


Interest revenue:
Interest revenue on $200,000 – June 30, 2018. On March 31, 2019 note was discounted at the Bank of Commerce.
200000X6%X3/12 = $3000
To determine interest rate on $60,000:
Interest on $200000 X 6% X 6/12 = $6000
Interest receivable $6800
Interest on $60000 note= 6800-6000 = $800 (for Nov and Dec)
Interest per month $800/2=$400
Annual interest 400X12 =$4800
60000 X rate % = 4800
Rate% = 4800/60000
Rate%= .08 or 8%
Interest revenue on $60,000 – Oct 31,2018
60000X8%X10/12= $4000
Total interest revenue = 3000 + 4000 = $7000
Bad debt expense
Beginning Account receivable (218000+24000) $             242,000
Sales $         1,340,000
Less – Cash collected $       (1,280,000)
Less – Write off $             (22,000)
Ending Account receivable $             280,000
It is stated that bad debt expense is adjusted at 10% of ending account receivable. 10% of $280000 = $28000
Allowance for uncollectible accounts:
Pre-adjustment balance- Beg. allowance 24000
Add: bad debt expense $$$
Less: write off -22000
Post adjustment balance – ending allowance 28000
24000 + $$$ - 22000 =28000
$$$ = 28000+22000-24000
Bad debt expense = $26000
Loss on note receivable:
Interest on $200000 note from 6/30/2018 – 3/31/2019
200000X6%X9/12= $9000
Cash proceeds received from discounted note
Face value of note $             200,000
Interest of 6% of maturity: 200000 X6% $               12,000
Total maturity value $             212,000
Less discount: 212000 X 8% X 3/12 $               (4,240)
Total cash proceeds $             207,760
Carrying value of note: 200000 + 9000 = $209000
Net loss on sale of note receivable: 209000 – 207760 = $1240